OpenStudy (anonymous):

how to get the derivative of 2^sin(x)

OpenStudy (anonymous):

I am assuming you set it to y, and derive by taking the natural log to bring sinx down?

OpenStudy (anonymous):

But I'm sure there is an easier way...

OpenStudy (zarkon):

\[\frac{d}{dx}\left[a^{f(x)}\right]=a^{f(x)}f'(x)\ln(a)\]

OpenStudy (anonymous):

So, (2^sinx)'=(2^sinx)(-cosx)ln2 ...?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Lovely! Thank you very much.

OpenStudy (zarkon):

no...cos(x)...not -cos(x)

OpenStudy (anonymous):

OH hahaha my bad -_-' Thank you for the correction.

OpenStudy (zarkon):

\[\frac{d}{dx}\sin(x)=cos(x)\]

OpenStudy (anonymous):

Right. I knew that. :S Thanks! :D

OpenStudy (anonymous):

Good catch by the way.

OpenStudy (zarkon):

That's my job ;)